### v2

# Integrating Sequential and Relational Modeling for User Events: Datasets and Prediction Tasks

Abstract
User event modeling plays a central role in many machine learning applications, with use cases spanning e-commerce, social media, finance, cybersecurity, and other domains. User events can be broadly categorized into personal events, which involve individual actions, and relational events, which involve interactions between two users. These two types of events are typically modeled separately, using sequence-based methods for personal events and graph-based methods for relational events. Despite the need to capture both event types in real-world systems, prior work has rarely considered them together. This is often due to the convenient simplification that user behavior can be adequately represented by a single formalization, either as a sequence or a graph. To address this gap, there is a need for public datasets and prediction tasks that explicitly incorporate both personal and relational events. In this work, we introduce a collection of such datasets, propose a unified formalization, and empirically show that models benefit from incorporating both event types. Our results also indicate that current methods leave notable room for improvements. We release these resources [^1] Link to the datasets and prediction tasks. Link to the dataset construction and experimentation code. to support further research in unified user event modeling and encourage progress in this direction.
1 Introduction
Modeling user events is a central task in machine learning with broad applications across various domains [purificato2024user, martin2021survey, deng2024advances]. In e-commerce, it is used to capture user preferences for personalized ranking and product recommendation [he2023survey, zeng2019user]. In social media platforms, event modeling supports feed optimization and engagement prediction by inferring user interests over time [abdelhafez2013survey, piao2018inferring, fang2023snapchat]. Financial systems leverage user behavior data for fraud detection, credit risk assessment, and behavioral profiling [hernandez2024fraud, mashrur2020risk, ozbayoglu2020deep, hojaji2022behavioral]. Online services such as search and streaming platforms rely on user event sequences for content recommendation under real-time constraints [zhao2023mbsrs, gayoavello2009session, covington2016youtube, amatriain2013netflix]. In cybersecurity, modeling user and system events is essential for detecting anomalies and preventing intrusions [lashkari2017survey, wang2017clickstream]. These applications demonstrate the importance of building models that can effectively capture complex, context-dependent user behavior from event sequences.
User events can be broadly categorized into personal and relational events. Personal events involve only a single user and reflect individual actions, such as searching for content, viewing items, or posting updates. In contrast, relational events involve interactions between two or more users, such as following another user, co-editing a document, or exchanging messages. Traditionally, these two types of events are often modeled separately. Relational events are commonly modeled using graph-based approaches that capture structural dependencies and interaction patterns among users [gao2023survey, guo2020survey, wu2022graph, li2024graph]. On the other hand, personal events are typically modeled as sequences using recurrent or attention-based architectures to capture temporal dependencies in personal event histories [Chen2024ASO, pan2024survey, boka2024survey, sun2019bert4rec, Tan2016ImprovedRN, Kang2018SelfAttentiveSR, liguori2023modeling].
Figure 1: An illustration of personal and relational events in e-commerce. Personal events involve a single user, such as login, search, view, or purchase. Relational events involve interaction between two users, such as sending a gift or commenting on another user’s review.
There have been efforts in the graph area to capture both structural and temporal dependencies using temporal graph formalizations (such as CTDG [nguyen2018continuous, tgn_icml_grl2020, kazemi2020representation]) and models built on top of these formalizations (such as TGAT [xu2020inductive], TGN [tgn_icml_grl2020], DyRep [trivedi2019dyrep], TNCN [zhang2406efficient], and others [luo2022neighborhood, yu2023towards, gravina2024long]). However, these approaches primarily focus on the temporal dependencies of relational events while neglecting personal events. For example, the formalization used in the Temporal Graph Benchmark (TGB) papers [huang2023temporal, gastinger2024tgb] and recent temporal graph models [zhang2406efficient, luo2022neighborhood, yu2023towards, gravina2024long] defines a temporal graph as a stream of triplets consisting of source, destination, and timestamp. Personal events that involve only a single entity cannot be directly represented under this formulation. One workaround is to convert all personal events into nodes and define personal events as triplets of user node, event node, and timestamp. However, this construction is not as straightforward for capturing temporal dependencies in personal event histories compared to sequence-based modeling.
Going back to the personal and relational event category, in many application domains, the number of personal events is typically much larger than that of relational events. For example, in e-commerce platforms, as illustrated in Figure 1, users often view products, search for items, or add products to their cart, whereas relational interactions, such as referrals, sending gifts, or socially engaged reviews, are less frequent. In financial systems, customers routinely perform account queries, check balances, or initiate transactions, while peer-to-peer interactions such as money transfers or joint account actions are relatively infrequent. In cybersecurity systems, personal events may include actions like logging in, accessing files, or executing processes, while relational events, such as remote connections to other users, or file sharing between users, occur less frequently. Despite their higher volume, personal events are often underrepresented in existing graph-based formulations, which tend to prioritize relational structure. In practice, however, both personal and relational events carry complementary signals, and many predictive tasks, such as item recommendation, fraud detection, customer profiling, and behavior forecasting, benefit from capturing both types of information.
Even though there is a need to capture both personal and relational events in many application domains, prior work has rarely considered them together. Practitioners often simplify the complexity of user event modeling by adopting either a graph or a sequence formalization, as most machine learning models are developed within one of these frameworks. As a result, one type of event, typically the less convenient to represent, is often ignored entirely, leading to an incomplete view of user behavior. To build a more comprehensive understanding of user event modeling, there is a need for public datasets and benchmark tasks that explicitly incorporate both event types. Such resources would provide a foundation for developing and evaluating models that integrate these complementary signals.
Summary of Contributions.
In this work, we aim to support the study of user event modeling that incorporates both personal and relational events. Our contributions are as follows:
  • We curate, pre-process, and release a collection of public datasets and prediction tasks that explicitly include both personal and relational events.
  • We introduce a new formalization for user event modeling that captures both personal and relational events.
  • We empirically demonstrate that incorporating both personal and relational events improves performance on a range of prediction tasks.
  • We show that existing models, originally developed for either sequential or relational data only, are less well suited for this event modeling setting, leaving room for future improvements.
  • We invite the community to use these resources and close the gap in unified user event modeling.
2 Related Works
We discuss the most relevant works in this section, while a more comprehensive list of the related works can be seen in Appendix D.
Event sequence modeling, such as temporal point processes [mei2017neural, zuo2020transformer, zhang2020self], deals with event stream data $(e_{1},t_{1}),(e_{2},t_{2}),\cdots$, where each $e_{i}$ is an event type drawn from an event set $\mathcal{E}=\{1,2,\cdots,|\mathcal{E}|\}$, and $0\leq t_{1}\leq t_{2}\leq\cdots$ are the times of occurrence. In user event sequence modeling, such as personalized event prediction tasks [letham2013sequential, boyd2020user, yang2022towards], we have $N$ users, denoted as $\mathcal{U}=\{u_{1},u_{2},\cdots,u_{N}\}$. Each user has their own sequence of events, denoted $\text{Seq}(u)=[(e_{1},t_{1})^{(u)},(e_{2},t_{2})^{(u)},\cdots]$. In some applications, exact timestamps are replaced with discrete time steps, simplifying the definition to $\text{Seq}(u)=[e_{1}^{(u)},e_{2}^{(u)},\cdots]$. In other variants of sequence modeling, the event space may be restricted to a single domain, for example, products consumed by a user in sequential recommendation tasks [hidasi2015session, kang2018self, sun2019bert4rec]. While these user event sequence models capture the complexity of event sequences within a user, they lack the ability to encode user-to-user interactions.
Graph modeling, on the other hand, explicitly encodes user-to-user interactions through its node and edge abstraction, $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ denote the node and edge sets, respectively. Temporal graph abstractions, such as CTDG [nguyen2018continuous, tgn_icml_grl2020], incorporate temporal dynamics by representing a graph as a sequence of time-stamped events $\mathcal{G}=\{x(t_{1}),x(t_{2}),\cdots\}$. Each event $x(t)$ can be either (1) a node-wise event (e.g., node addition or feature update) represented by a feature vector $\mathbf{\bm{v}}_{i}(t)$, where $i$ is the node index, or (2) an interaction event between node $i$ and node $j$ represented by a temporal edge $\mathbf{\bm{e}}_{i,j}(t)$. Another CTDG formulation by kazemi2020representation describes a temporal graph as a pair consisting of an initial graph and an observation set $(\mathcal{G},\mathcal{O})$. The observation set describes the evolution of the graph, with each observation is represented as a tuple (event type, event, timestamp), where the event type may include edge addition, edge deletion, node addition, node deletion, node splitting, node merging, etc. In practice, however, many temporal graph models [zhang2406efficient, luo2022neighborhood, yu2023towards, gravina2024long] omit node-level events altogether and instead define a temporal graph as a sequence of interaction events expressed as triplets, $\mathcal{G}=\{(u_{1},v_{1},t_{1}),(u_{2},v_{2},t_{2}),\cdots\}$, where $u$ and $v$ represent the source and destination nodes.
3 Problem Formalization
Notations.
For our Personal and Relational User Event Sequence (PRES) modeling, we take inspiration from the standard user event sequence modeling. We denote a set of users (which can also be a customer, account, entity, etc.) as $\mathcal{U}=\{u_{1},u_{2},\cdots,u_{N}\}$, where $N$ is the number of users. Each user has their own sequence of events that occur over time. For example, the sequence for user $u_{i}$ is denoted as $\text{Seq}(u_{i})=[(e_{1},t_{1}),(e_{2},t_{2}),\cdots,(e_{M_{i}},t_{M_{i}})]$, where $e$ describes an event, $t$ describes the time at which the event occurs, and $M_{i}$ denotes the number of events for user $u_{i}$. Each user may have a different number of events in their event sequence. We denote the set of all user sequences by ${\mathcal{S}}=\left\{\text{Seq}(u)\mid u\in\mathcal{U}\right\}$. An event may come from two different event sets: the personal event set and the relational event set. The personal event set contains a set of events that can occur for an individual user; $p\in\mathcal{P}\triangleq\{1,2,\cdots,|\mathcal{P}|\}$. The relational event set contains a set of all possible events $r\in\mathcal{R}\triangleq\{1,2,\cdots,|\mathcal{R}|\}$, which involve a relation from one user to another. Thus, an event can be defined by a personal event $e=p$, or a relational event tuple $e=(r,v)$, where $v$ is another user.
Difference from temporal graph formulation.
Although motivated by a similar goal of integrating temporal and relational information, our formulation differs from temporal graphs in several ways:
  • In the temporal graph formulation, events are ordered globally across all users, similar to the standard (non-user-based) event sequence modeling. In contrast, PRES follows personalized user-event sequence modeling, where each user has their own separate sequence of events.
  • The temporal graph formulation places greater emphasis on modeling interaction events between two nodes, often neglecting user events that do not involve another user. Many recent temporal graph models, such as TNCN [zhang2406efficient], NAT [luo2022neighborhood], DyGFormer [yu2023towards], and others [gravina2024long, lu2024improving, gao2025hyperevent]; recent temporal graph benchmarks, such as TGB [huang2023temporal] and TGB 2.0 [gastinger2024tgb]; as well as popular graph frameworks such as PyTorch Geometric [fey2019fast], even reduce the formulation to a sequence of triplets representing interaction events with no support for encoding personal events. By contrast, PRES treats personal events as first-class components. This is particularly relevant in many areas mentioned in the introduction, where the number of personal events is far larger than that of relational events.
  • Even for the temporal graph formulation that accommodate node-wise events, it does so by updating node feature vectors. This differs from event sequence modeling (including PRES), where each event is drawn from a discrete set of events, a property that may not be easily encoded as feature changes. To represent discrete personal events, temporal graph models may convert the events into nodes and then build a heterogeneous dynamic graph that encodes the event as an interaction between a user node and this newly created event node. In some events, such as ‘viewing product’, this approach works fine. However, it does not work well for representing personal events that describe intransitive actions (e.g., ‘sign in’, ‘login’, or ‘subscribe’) or status changes (e.g., ‘payment successful’, ‘login unsuccessful’, or ’request denied’).
  • In graph models, every entity must be a node. This becomes problematic when the entity has a hierarchical structure. For example, in a location-based social network application, when modeling a user’s location check-in event, one may want to capture hierarchical information such as continent, country, state, city, neighborhood, and block. Representing the location check-in as a single node in a graph model loses this hierarchical structure. In contrast, PRES allows more flexibility: using a sequence model, events can be freely tokenized. A check-in event can be split into multiple tokens, each corresponding to a different level of location granularity.
When we compare with the CTDG formulation by kazemi2020representation, all the concerns above remain. While this formulation allows more flexible observations, it is still restricted to events that describe the evolution of the graph through changes in nodes or edges, such as edge additions, edge deletions, node additions, node deletions, node splitting, or node merging.
4 Datasets and Prediction Tasks
4.1 Dataset Information
Table 1: Dataset Statistics
Properties brightkite gowalla amazon-clothing amazon-electronics github
Personal Events check-in check-in product rating product rating github activity
Relational Events friendship friendship co-review co-review collaboration
# Users 58,228 196,591 185,986 254,064 3,669,079
# Events 5,130,866 8,342,943 1,591,947 2,938,178 102,878,895
# Personal Events 4,702,710 6,442,289 1,573,869 2,281,128 95,974,149
# Relational Events 428,156 1,900,654 18,078 657,050 6,904,746
# Unique Events 628,519 1,169,154 846,052 529,198 24
# Unique Timestamps 4,506,822 5,561,957 3,464 5,373 2,675,990
# Users w. pers. events 51,406 107,092 185,986 254,064 3,669,079
# Users w. rel. events 58,228 196,591 5,017 49,852 441,958
# Users w. both events 51,406 107,092 5,017 49,852 441,958
We curated user event datasets from multiple domains [^2] The datasets presented in this paper are used solely to demonstrate the claims made in this work and are not reflective of any datasets leveraged by Capital One., following recent graph dataset curation works [huang2023temporal, gastinger2024tgb, yi2025tgb, hu2020open, hu2021ogb], and then processed them according to our formalization in Section 3. The data is stored in CSV format with the columns: uid, timestamp, event_set, event, and other_uid (See Appendix B for details). The uid is a numerical user ID, whereas event_set indicates whether the event is personal or relational. For relational events, other_uid refers to the other user involved in the relation; for personal events, this column is null. The statistics of each datasets is shown in Table 1. More details on collection and dataset license are available in Appendix A. The data processing code are open-sourced and descriptions are available in Appendix C.
pres-brightkite. This dataset contains location check-ins and friendship history of Brightkite users, a location-based social networking platform. It was originally collected by cho2011friendship and published in the SNAP Dataset Repository [snapnets]. Personal events consist of sequences of location check-ins. We convert the original latitude and longitude coordinates into Geohash-8 representations [niemeyer2008geohash, morton1966computer], short alphanumeric strings encoding geographic locations. Nearby locations share similar geohash prefixes, while distant ones differ. Example geohashes include 9v6kpmr1, gcpwkeq6, and u0yhxgm1. Relational events capture friendship connections among users. The dataset includes 58,228 users and 5,130,866 events. In this dataset, only the personal events include timestamp information; relational events do not have associated timestamps.
pres-gowalla. The dataset also contains the location check-in and friendship history of another social network platform, Gowalla. It was also originally collected by cho2011friendship and published in the SNAP Repository [snapnets]. We processed and formatted the data following the same approach used for pres-brightkite. The dataset contains personal events from geohash check-ins and relational events from friendship connections, totaling 8,342,943 events from 196,591 users.
pres-amazon-clothing. The dataset contains Amazon product reviews and ratings in the Clothing, Shoes and Jewelry category, spanning from May 1996 to July 2014. The raw data was originally collected by mcauley2015image. In this dataset, we define personal events as sequences of product IDs and ratings reviewed by a user, for example: B [^0] MLDCZ2:5 and B001OE3F08:3. Relational events represent co-review patterns, where two users have co-reviewed at least three products. The dataset contains event sequences from 185,986 users, with a total of 1,591,947 events.
pres-amazon-electronics. The dataset contains Amazon product reviews and ratings in the Electronics category, originally collected by mcauley2015image. As in pres-amazon-clothing, personal events are defined as sequences of product IDs and ratings, while relational events capture co-review patterns. In total, the dataset contains 2,938,178 events from 254,064 users.
pres-github. This dataset contains GitHub user activity from January 2025, collected from the GH Archive. Personal events include actions such as Push, CreateBranch, CreateRepository, PullRequestOpened, IssuesOpened, and Fork. Relational events represent project collaboration, where two users are linked if both contributed at least five commits or pull requests to the same repository. The dataset includes 102,878,895 events from 3,669,079 users. In this dataset, only personal events include timestamp information; relational events do not have associated timestamps, similar to pres-brightkite and pres-gowalla.
Variability of the datasets.
As shown in Table 1, the PRES datasets vary significantly across multiple aspects. The number of users ranges from around 58 thousand in pres-brightkite to more than 3.5 million in pres-github. The number of events also varies, from approximately 1.5 million in pres-brightkite to over 100 million in pres-github. The ratio between relational and personal events ranges from around 1:3 in pres-gowalla to approximately 1:80 in pres-amazon-clothing. The number of unique events also differs widely, from just 24 in pres-github to more than 1 million in pres-gowalla. In addition, we observe variability in the number of users having personal events, relational events, and both. Some datasets have more users with relational events than with personal events (e.g., pres-brightkite, pres-gowalla), while others show the opposite trend (e.g., pres-amazon-clothing, pres-amazon-electronics, pres-github). These differences in dataset properties present distinct challenges for modeling user events in each dataset.
4.2 Prediction Tasks
From the PRES datasets, we define two prediction tasks: one for relational events and one for personal events. These tasks are designed to enable fair comparisons between graph-based, sequence-based, and hybrid models. Relational event prediction focuses on predicting future or held-out subset of user-to-user interactions, similar to link prediction. Personal event prediction aims to predict the likelihood of future occurrence of personal events without requiring exact timestamps, for example, predicting the next 20 personal events given a user’s first 100. In both tasks, observed events are compared against negative samples drawn from events not associated with the user. For reproducibility, pre-generated negative samples for validation and test sets are provided in the dataset repository.
Relational event prediction tasks.
The corresponding tasks for pres-brightkite and pres-gowalla involve friend recommendation. We construct the training data by randomly splitting all relational events into 70% training, 10% validation, and 20% test sets. We also generate negative samples for the validation and test sets. Following gastinger2024tgb, we adopt a 1-vs-1000 negative sampling scheme, in which 1,000 negative events are sampled for each relational event in the prediction set. Negative samples are drawn via uniform random sampling of users, excluding those who already have relational events with the target user in the training set.
For the pres-github dataset, the relational event prediction task is defined as collaboration prediction, which involves predicting which users collaborate with a given user. The train, validation, and test splits follow the same procedure as in pres-brightkite, including the sampling method. However, due to the large size of the dataset, we adopt a 1-vs-300 negative sampling scheme.
For the pres-amazon-clothing and pres-amazon-electronics datasets, the task is predicting co-review relationships, i.e., which users share at least three products they reviewed. Co-review patterns can reveal how one account may be related to another, which in some cases can help detect fraudulent review syndicates. In these datasets, relational events have timestamp information, i.e., the first time the co-review condition is met. As such, the train, validation, and test splits respect event timestamps. Specifically, we split each user’s relational events by taking the last 20% for test, the previous 10% for validation, and the rest for training. To manage large histories of some users, we cap test and val sets at 20 and 10 events per user, respectively. Personal events are also split into ‘observed’ and ‘unobserved’ sets based on the timestamp cut-off in the relational event split, with only the observed set used for training. As in pres-brightkite, we adopt a 1-vs-1000 negative sampling scheme.
Personal event prediction tasks.
The task for pres-brightkite and pres-gowalla is to predict the likelihood of a user checking in at a given geohash location in the future. We split each user’s personal events by taking the last 20% for test, the previous 10% for validation, and the rest for training. We also cap the number of events in the test and validation sets to at most 20 and 10 per user, respectively. Since personal events are more frequent than relational ones, we adopt a 1-vs-500 negative sampling scheme. As geohash strings encode hierarchical spatial information (e.g., earlier characters represent broader regions), we apply stratified hierarchical sampling. Specifically, negatives are stratified by shared geohash prefixes, from matching the first five characters to none, ensuring a mix of nearby and distant locations.
For the pres-amazon-clothing and pres-amazon-electronics datasets, the task is to predict future products a user will review and the corresponding ratings, as denoted in their personal event data. We adopt the same train/val/test split strategy as in pres-brightkite, along with a 1-vs-500 negative sampling scheme. Negative samples for each personal event (e.g., B001OE3F08:3) are drawn from three sources: (1) the same product with different ratings (e.g., B001OE3F08:1, B001OE3F08:5); (2) other personal events not in the user’s training data; and (3) samples from the second set with randomly perturbed ratings.
In the pres-github dataset, the number of unique events in the personal event set is only 24, corresponding to the list of possible GitHub activities. Thus, the task construction used in the previous datasets is not applicable to pres-github. We decided to omit this dataset from the set of datasets used for creating personal event prediction tasks.
Full event sequence.
In addition to the datasets containing prediction tasks described above, we also publish a version of each dataset that includes all personal and relational events for all users, without any assigned tasks, train/val/test splits, or pre-specified negative samples. This is intended to facilitate future works that may wish to generate other prediction tasks not covered in this paper.
5 Experiments
5.1 Relational event prediction tasks
Table 2: Performance results for relational event prediction tasks (all metrics are in percent).
Method pres-brightkite pres-gowalla
Metric MRR Hits@5 Hits@10 Hits@50 Hits@100 MRR Hits@5 Hits@10 Hits@50 Hits@100
Static graph models on relational event graph
[…]
Static graph models on relational event graph + sequence embedding from personal event data
[…]
Static graph models on relational event graph + personal event nodes
[…]
Temporal graph models on relational event graph + personal event nodes
[…]
Method pres-amazon-clothing pres-amazon-electronics
Metric MRR Hits@5 Hits@10 Hits@50 Hits@100 MRR Hits@5 Hits@10 Hits@50 Hits@100
Static graph models on relational event graph
[…]
Static graph models on relational event graph + sequence embedding from personal event data
[…]
Static graph models on relational event graph + personal event nodes
[…]
Temporal graph models on relational event graph + personal event nodes
[…]
Experiment setup. We perform relational event prediction experiments on all five PRES datasets, following the task setup described earlier. We evaluate several sets of baseline methods:
  1. 1.
    In the first set, we use only relational event data. We construct a user graph where edges represent relational events between two users, ignoring timestamp information. We then run static graph methods, GCN [kipf2017semi], GAT [velivckovic2018graph], and Graph Transformer (TConv) [shi2021masked] on this graph.
  2. 2.
    In the second set, we use a biderectional transformer sequence model (BERT) [devlin2019bert], to encode each user’s last 100 personal events from the training set. The resulting user embedding is added as input to the GCN, GAT, and TConv models from the first set, denoted as GCN+S, GAT+S, and TConv+S, respectively.
  3. 3.
    In the third set, we convert each unique personal event into a node and add it to the user graph from the first set, creating edges between users and their personal event nodes. As in the second set, we use only the last 100 personal events per user. We then run GCN, GAT, and TConv on this graph, denoted as GCN_RP, GAT_RP and TConv_RP.
  4. 4.
    Lastly, based on the graph containing user and personal event nodes from the third set, we add timestamp information to construct a temporal graph. For datasets that lack timestamps for relational events, we inject these events randomly into the sequence of personal events. We then run temporal graph models, TGN [tgn_icml_grl2020], DyRep [trivedi2019dyrep], and TNCN [zhang2406efficient] on this graph.
The sequence model for capturing personal events in the second set is designed as a masked token prediction task using a BERT model with a masking probability of 0.3. A key benefit of using transformer-based models is flexibility in event tokenization. In pres-brightkite and pres-gowalla, personal events are 8-character geohash strings (e.g., 9q8yyk8y|9q8vzj5b|9q8vyzwk). Since geohashes encode hierarchical geographic information, we apply hierarchical tokenization by splitting each into four two-character tokens with added prefixes (e.g., gh12-9q, gh34-8y, gh12-yk, gh12-8y). This roughly mimics hierarchical location modeling, such as identifying continent, country, city, and neighborhood. For the pres-amazon datasets, we apply similar tokenization by splitting each event into three product tokens and one rating token. We do not apply token splitting for pres-github.
For performance evaluation, following prior benchmarks [huang2023temporal, gastinger2024tgb, yi2025tgb], we use ranking-based metrics: Mean Reciprocal Rank (MRR) and Hits@k, evaluated at various $k$ depending on the number of negative samples. Each baseline is run five times with different random seeds, and we report the mean and standard deviation of the results.
Experiment results. Table 2 and Table 3 show the experiment results, with additional results available in Appendix F. In each table, bold numbers indicate the best-performing model on a given metric, and underlined numbers indicate the second best. As each dataset has its own characteristics, the results vary across datasets. However, there are some emerging patterns in the results that we highlight below.
Table 3: Relational event predictions on pres-github.
Method MRR Hits@3 Hits@5 Hits@10 Hits@30
[…]
TGN Out of GPU Memory
DyRep Out of GPU Memory
TNCN Out of GPU Memory
  • The graph models with personal event sequence embeddings—GCN+S, GAT+S, and especially TConv+S—consistently perform well across all datasets and metrics. On pres-brightkite and pres-gowalla, TConv+S outperforms all models across all metrics, often by a wide margin, especially compared to methods that encode personal events as nodes. On pres-github, GAT+S shines, outperforming all methods on all metrics.
  • For the amazon datasets, TConv+S remains dominant on Hits@k metrics with larger $k$, achieving the best results at Hits@10, Hits@50, and Hits@100 on pres-amazon-clothing, as well as Hits@50 and Hits@100 on pres-amazon-electronics, while holding second place at other $k$ values. At these lower $k$, where TConv+S ranks second, the best model is TNCN, a temporal graph method. TNCN, however, does not perform well on Hits@k metrics with larger $k$.
  • With the exception of TNCN on the amazon datasets with small $k$ metrics, the performance of temporal graph methods (TGN, DyRep, and TNCN) using graphs with personal event nodes is noticeably lower compared to static graph models with sequence embeddings. For the large pres-github dataset, they encounter GPU out-of-memory errors, even when using small batch sizes.
  • In nearly all cases, converting personal events into nodes and adding them to the relational event graph is less effective than modeling personal events with a sequence model and using the user’s sequence embedding as an additional node features to the relational event graph. The performance of the “_RP” versions of static graph models is lower than that of the corresponding “+S” versions across nearly all datasets and metrics, with a few exceptions on the amazon datasets for Hits@k metrics with lower $k$.
Even though GCN+S, GAT+S, and TConv+S perform relational event prediction in two stages, first generating user embeddings from personal event sequences and then incorporating them into the graph learning process, they still perform well across datasets. In contrast, TGN, DyRep, and TNCN use a single-step approach that directly integrates temporal dynamics but operates on graphs where personal events are represented as nodes. These differences highlight an opportunity for future exploration of how best to represent the temporal dynamics of personal events within a user while jointly modeling the full structure that includes user-to-user relational events, in an end-to-end fashion.
5.2 Personal event prediction tasks
Table 4: Performance results for personal event prediction tasks (all metrics are in percent).
Method pres-brightkite pres-gowalla
Metric MRR Hits@3 Hits@5 Hits@10 Hits@50 MRR Hits@3 Hits@5 Hits@10 Hits@50
Sequential models
[…]
Graph models on personal event only graph
[…]
Graph models on personal and relational event graph
[…]
Method pres-amazon-clothing pres-amazon-electronics
Metric MRR Hits@3 Hits@5 Hits@10 Hits@50 MRR Hits@3 Hits@5 Hits@10 Hits@50
Sequential models
[…]
Graph models on personal event only graph
[…]
Graph models on personal and relational event graph
[…]
Experiment setup. We perform personal event prediction experiments on all PRES datasets except pres-github. In these experiments, we evaluate several sets of baseline methods:
  1. 1.
    The first model is a sequential model that uses only personal event data. We use a bidirectional transformer architecture (BERT) with a prediction head to compute the likelihood of a user having a particular personal event in the future. For each user, we use the last 100 personal events in the training set to predict the likelihood of future events.
  2. 2.
    In the second set, we use node2vec [grover2016node2vec] to learn the graph structure of relational events and generate a graph embedding for each user. We then incorporate the embedding into the BERT sequence model. We evaluate two versions of the model: (a) incorporating the graph embedding post transformer module and before the prediction head (BERT-n2v-p), and (b) using the embedding as a special input token to the transformer module (BERT-n2v-i).
  3. 3.
    In the third set, we use graph-based models on personal event–only data by creating a bipartite graph of user nodes and personal event nodes, based on the last 100 personal events per user. We run both static graph models (GCN and GAT) and temporal graph models (TGN and DyRep) on this graph.
  4. 4.
    In the last set, we augment the graph in the third set with relational event data by adding relational event edges between users. We then run GCN, GAT, TGN, and DyRep on this graph, denoted as GCN_PR, GAT_PR, TGN_PR, and DyRep_PR, respectively.
Similar to the sequence embedding used in relational event prediction tasks, we apply split tokenization for the BERT model in personal event prediction to allow more flexibility in modeling events. We use the same tokenization scheme for each dataset as described earlier. For evaluation, we report MRR and Hits@k at various values of $k$. Each baseline is run five times with different random seeds, and we report the mean and standard deviation of the results.
Experiment results. Table 4 shows the results for the personal event prediction task. As in the relational event task, results vary across datasets due to their unique characteristics, with even more variations in this setting. We discuss some of the results as follows.
  • The sequence models perform well on pres-brightkite across all metrics. The base BERT model, which uses only personal event data, already shows strong performance. Adding relational event node2vec embeddings may either improve or degrade performance. In pres-brightkite, adding the embedding after the transformer module reduces performance, while using it as a special input token improves it. However, the changes are relatively minor but sufficient to make BERT-n2v-i the best-performing model on pres-brightkite. Similar minimal changes are observed in other datasets.
  • The static graph model, GCN in particular, performs surprisingly well on pres-gowalla. The best performance is achieved by the GCN_PR model, which is trained on data containing both personal and relational events in a graph with user nodes and personal event nodes. GCN_PR also performs relatively well on pres-amazon-clothing and pres-amazon-electronics. However, the GAT-based models perform noticeably worse than their GCN counterparts.
  • The temporal graph models perform relatively well on the pres-amazon-clothing and pres-amazon-electronics datasets, particularly on the Hits@5 and Hits@10 metrics. TGN and DyRep perform better on graphs that include both personal and relational events. A notable exception is the Hits@50 metric.
The results show that there is no single model that consistently performs best across all datasets. Some models work well on certain datasets but not on others. The only consistent pattern is that the best-performing models usually use both personal and relational events. This opens up opportunities for designing better models that can effectively integrate both types of information.
5.3 Hypothesis on native end-to-end model architectures for PRES datasets
In the experiment results described above, we have demonstrated that current models leave notable room for performance improvements in both personal and relational prediction tasks. We hypothesize that a significant performance improvement can be achieved by end-to-end models trained simultaneously on both personal and relational event data. There are many ways to design such models; below are some examples:
  1. 1.
    A possible strategy is to start with a user-level Transformer architecture that takes the user event sequence as input, combined with a message-passing architecture that enables information transfer between two users connected by a relational event and propagates that information across the neighbors via a Graph Neural Network (GNN)-like architecture. This design capitalizes on the strengths of both the Transformer and GNN architectures. Transformer architectures have been shown to be the best model for handling sequential information, whereas GNN architectures excel in transferring information across connected users.
  2. 2.
    Another possible architecture candidate is to also start from the Transformer architecture for the sequence, where each sample represents a user event sequence, and allow the attention mechanism to propagate toward other user sequences that are connected via a relational event, in addition to the attention mechanism inside each user/sequence.
There could also be multiple other promising architectural paradigms for developing a fully end-to-end model. We leave the exploration of these alternative approaches for future work.
6 Conclusions and Limitations
In this work, we aim to advance user event modeling by introducing a unified framework that captures both personal and relational events. We curate and release a collection of public datasets with corresponding prediction tasks, all aligned under a formalization that integrates both event types to provide a more complete view of user behavior. Through empirical evaluation, we demonstrate that models leveraging both event types consistently outperform those using only one. We also show that existing methods, originally developed for either sequential or relational data, even with some adaptations to handle both (e.g., temporal graph models), are less effective across many of our prediction tasks. These findings highlight the need for further study of unified user event modeling.
A key challenge in this work is in the dataset curation process, as many public datasets have already been collapsed into either graph-only or sequence-only formats, often discarding personal or relational events in the process. While we were able to gather and unify a set of datasets that include both event types, they may not fully capture the diversity and complexity of user event modeling across domains. Another limitation is that our current formulation does not support event-level or user-level features, presenting an opportunity for future work to extend the framework toward feature-aware modeling.
Appendix A Dataset Documentation
All datasets presented in this paper are intended for academic research purposes, and their corresponding licenses are listed in this section. They are constructed from publicly available resources described below. In all cases, we perform anonymization by removing any personally identifiable information. User IDs in the original data are replaced with auto-incremented ID numbers.
Download links.
The datasets and tasks described in this paper are available for download from the following links:
Dataset source and license information.
Below, we describe how the source data was obtained and provide license information for each dataset:
  • pres-github. This dataset is based on GitHub data collected from the GH Archive website (https://www.gharchive.org/) using its HTTP JSON download link. It contains GitHub user activity from January 2025, and user IDs have been anonymized. Content from GH Archive is released under the CC-BY-4.0 license, while the associated code is released under the MIT license.
  • pres-amazon-clothing and pres-amazon-electronics. These datasets contain Amazon product reviews and ratings in their respective categories. Both are based on Amazon review data collected by mcauley2015image and hosted at: https://cseweb.ucsd.edu/˜jmcauley/datasets/amazon/links.html. The Amazon review content is licensed under the Amazon license:
    By accessing the Amazon Customer Reviews Library ("Reviews Library"), you agree that the Reviews Library is an Amazon Service subject to the Amazon.com Conditions of Use and you agree to be bound by them, with the following additional conditions:
    In addition to the license rights granted under the Conditions of Use, Amazon or its content providers grant you a limited, non-exclusive, non-transferable, non-sublicensable, revocable license to access and use the Reviews Library for purposes of academic research. You may not resell, republish, or make any commercial use of the Reviews Library or its contents, including use of the Reviews Library for commercial research, such as research related to a funding or consultancy contract, internship, or other relationship in which the results are provided for a fee or delivered to a for-profit organization. You may not (a) link or associate content in the Reviews Library with any personal information (including Amazon customer accounts), or (b) attempt to determine the identity of the author of any content in the Reviews Library. If you violate any of the foregoing conditions, your license to access and use the Reviews Library will automatically terminate without prejudice to any of the other rights or remedies Amazon may have.
  • pres-gowalla. This dataset contains user activity on the (now defunct) social network Gowalla. It was originally collected by cho2011friendship using the platform’s public API and published in the SNAP Dataset Repository [snapnets] (https://snap.stanford.edu/data/loc-Gowalla.html). The curator confirmed that SNAP datasets are free to use, but no specific license information is available.
  • pres-brightkite. This dataset contains user activity on the (also now defunct) social network Brightkite. It was also originally collected by cho2011friendship using the platform’s public API and published in the SNAP Dataset Repository [snapnets] (https://snap.stanford.edu/data/loc-brightkite.html). The curator confirmed that SNAP datasets are free to use, but no specific license information is available.
Appendix B Dataset Contents
Examples of dataset contents.
To illustrate the structure of the curated datasets, we provide examples of user event sequences from several pres datasets. Each table includes both personal and relational events, showing how different types of user activity are represented in our format.
  • pres-brightkite and pres-gowalla
    Table 5: Example of user event sequence in pres-brightkite and pres-gowalla.
    uid timestamp event_set event other_uid
    39 1206596784 personal 9xj6hwkm <NA>
    39 1206596838 personal 9xj3fynm <NA>
    39 1206596871 personal 9xj3fynm <NA>
    39 1235862855 personal 9xj65423 <NA>
    39 1250883230 personal 9xj65423 <NA>
    39 1254535157 personal 9xj5skbn <NA>
    39 1254535193 personal 9xj5sm00 <NA>
    39 1283443369 personal 9q8yyyhs <NA>
    39 <NA> relational friendship 0
    39 <NA> relational friendship 30
    39 <NA> relational friendship 105
    39 <NA> relational friendship 1190
  • pres-amazon-clothing and pres-amazon-electronics
    Table 6: Example of user event sequence in pres-amazon-clothing and pres-amazon-electronics.
    uid timestamp event_set event other_uid
    254057 1375401600 personal B [^0] A6PPOK:3 <NA>
    254057 1377302400 personal B003TMPHOU:5 <NA>
    254057 1377302400 personal B004A81PJI:4 <NA>
    254057 1377302400 personal B0054R4AXW:5 <NA>
    254057 1377302400 personal B005CPGHAA:5 <NA>
    254057 1377302400 personal B007FNXMEQ:5 <NA>
    254057 1377302400 personal B007IV7KRU:5 <NA>
    254057 1377302400 personal B007WAWHD4:5 <NA>
    254057 1377302400 personal B008AST7R6:5 <NA>
    254057 1377302400 personal B008R56H4S:5 <NA>
    254057 1404086400 relational co-review_product 107741
  • pres-github
    Table 7: Example of user event sequence in pres-github.
    uid timestamp event_set event other_uid
    3669059 1738288160 personal PullRequestReviewCreated <NA>
    3669059 1738288191 personal PullRequestReviewCreated <NA>
    3669059 1738288198 personal PullRequestClosed <NA>
    3669059 1738288200 personal Push <NA>
    3669059 1738288206 personal PullRequestClosed <NA>
    3669059 1738288207 personal Push <NA>
    3669059 1738288217 personal DeleteBranch <NA>
    3669059 1738288219 personal DeleteBranch <NA>
    3669059 <NA> relational collaborate 824409
    3669059 <NA> relational collaborate 3126262
(a) pres-brightkite
(b) pres-gowalla
(c) pres-amazon-clothing
(d) pres-amazon-electronics
Figure 2: Histogram of the number of events per user in each dataset.
Figure 3: Histogram of the number of event per user in pres-github.
Event statistics.
To characterize user events, we include histograms in Figure 2 and Figure 3 showing the distribution of event counts per user in each dataset. These histograms are constructed by computing the number of events associated with each user and aggregating how many users fall into each count bucket. The y-axis is log-scaled to highlight the long-tailed nature of user behavior, where the majority of users generate only a small number of events, while a much smaller group contributes disproportionately large volumes of activity. This skew is common across datasets and presents both challenges and opportunities for modeling.
Appendix C Dataset Construction
This section provides more details about how all of the PRES datasets are created. In short, the data processing files are available in our GitHub repo, following the format create_datasets/process_X.ipynb for first generating the processed files, and create_datasets/task_X.py for subsequently generating the prediction tasks (including negative sampling for test set), where X is the dataset name.
C.1 pres-brightkite and pres-gowalla
The raw data of users’ friendships and check-in histories are stored in text files. We read the friendship files into a dataframe and reformat them into our standard format. The check-in history file contains the latitude and longitude of check-in locations. After reading the files, we convert the latitude-longitude encoding of locations into Geohash-8 string representations and convert the check-in times into Unix timestamp format. Relational events denote friendship between two users, and there are no timestamps associated with this friendship from the raw data. We combine both event types, sort the data by uid and timestamp, and save it into a CSV file (in our dataset repository on Kaggle, this is processed/X_all_events.csv).
Now that we have all events saved in a standardized format, we can create our splits and pre-generate negative samples for reproducibility within create_datasets/task_brightkite.py and create_datasets/task_gowalla.py. For both tasks, we begin by splitting all events into separate dataframes for personal and relational events.
Relational Task: Friendship Prediction
To generate the relational event prediction task, we perform a random split of 70%/10%/20% on the relational event dataframe to generate the train/eval/test datasets, and store them into CSV files. Following gastinger2024tgb, we adopt a 1-vs-1000 negative sampling scheme, in which 1,000 negative events are sampled for each relational event in the prediction set. Negative samples are drawn via uniform random sampling of users, excluding those who already have relational events with the target user in the training set.
Personal Task: Check-in Prediction
To generate the personal event prediction task, we perform timestamp-based splitting on the personal event dataframe into train/val/test sets. We split each user’s personal events by taking the last 20% for the test set, the previous 10% for validation, and the remaining 70% for training. We cap the number of events in the test and validation sets to at most 20 and 10 per user, respectively. Because personal events are more frequent than relational ones, we adopt a 1-vs-500 negative sampling scheme. As geohash strings encode hierarchical spatial information (e.g., earlier characters represent broader regions), we apply stratified hierarchical sampling. Specifically, negatives are stratified by shared geohash prefixes, from matching the first five characters to none. Varying the number of matching characters ensures our negatives contain a mix of nearby and distant locations.
C.2 pres-amazon-electronics and pres-amazon-clothing
The input zipped JSON containing the raw data is loaded into a Pandas DataFrame. Each row corresponds to a single review by a single user. We drop rows belonging to users with fewer than 5 product reviews, and create anonymous user IDs from the input ’reviewID’ column.
Personal events here are simply all the remaining rows; the ’event’ column is created by concatenating the product ID with the rating given to it in that review. Relational events are created by finding users who have co-reviewed at least 3 of the same products. Timestamps for personal events are naturally the time at which the review was posted. For relational events, the timestamp corresponds to when this co-review condition is first met.
The schema for both of these personal and event dataframes are homogenized. At the end of process_amazon-X.ipynb, we combine both the relational and personal event data and write it out as a single CSV. This is now the input into task_amazon.py for each dataset.
Both tasks use a 70% train, 10% validation and 20% test split paradigm, where the test events are the latest 20% of events (either personal or relational) per user, with the validation events immediately preceding the test events. To manage large histories of some users, we cap test and val sets at 20 and 10 events per user, respectively.
Relational Task: Co-Review Prediction
We apply the above logic to only the relational events, and can now identify the maximum timestamp (aka cut-off time) per user in the training data. This is used to split all personal events into ’observed’ (before cut-off time) and ’unobserved’ (after cut-off time). Only observed personal events are used for training. Mirroring the logic for pres-brightkite and pres-gowalla’s relational task, we adopt a 1-vs-1000 negative sampling scheme. Each negative corresponds to a uniform random sampling of user IDs that do not have a true co-review relationship with the user ID of interest.
Personal Task: Product Prediction
Similarly to the relational task, we apply the splitting logic only to personal events. We use the timestamp of the latest train event per user to identify which of that user’s relational events are included for training. We adopt a 1-vs-500 negative sampling scheme, aligning with pres-brightkite and pres-gowalla’s personal task logic. Negative samples for each personal event (e.g., B001OE3F08:3) are drawn from three sources: (1) the same product with different ratings (e.g., B001OE3F08:5); (2) other personal events not in the user’s training data; and (3) samples from the second set with randomly perturbed ratings.
C.3 pres-github
The raw data of GitHub users’ activities comes as a compressed JSON. We unpack this into a CSV in create_datasets/github_extract.py. We aim to remove bot accounts that could add substantial noise by dropping any rows where the user contains [bot] in their name or if that user conducted 100,000 or more actions in the dataset.
Next, in create_datasets/process_github.ipynb we additionally remove users with fewer than 3 actions. Relational events connect users who have pushed or opened pull requests at least 5 times in the same repository; predicting relational events in this context is about predicting collaborators. Personal events correspond to the rows in the original dataset, which represent only 24 unique GitHub activities. Thus, the personal event task construction used in the previous datasets is not applicable to pres-github. We decided to omit this dataset from the set of datasets used for creating personal event prediction tasks. These personal and relational events have their schemas homogenized and are concatenated. The result is saved as processed/github_all_events.csv on Kaggle datasets.
Finally, create_datasets/task_github.py is used to create train/validation/test splits and negatives for the relational event prediction task. The logic here follows pres-brightkite, including the split proportions and limitations on the maximum number of events withheld for validation and test. The notable difference is that we adopt a 1-vs-300 negative sampling scheme due to the large size of the dataset.
Appendix D Additional Related Works
D.1 Event sequence.
Event sequence modeling is a broad topic that covers many different domains which share a similar goal of understanding and potentially predicting future events given past history. In healthcare, being able to predicting patient’s upcoming medical visits enables proactive care and better resource allocation for healthcare providers. In manufacturing and industrial setting, modeling sequences of equipment sensor readings or machine states allow for early detection of faults, enabling predictive maintenance and reducing downtime. These problems share common challenges such as modeling temporal dependencies, handling irregular and asynchronous observations.
Temporal point processes (TPPs) provide a powerful tool for modeling discrete events as stochastic processes in continuous time. Classical models like the Poison process and the Hawkes process [hawkes1971spectra] allow for explicit modeling of event dynamics, including self-excitation and mutual inhibition, through parameterized intensity functions that govern the likelihood of future events. Recent work has introduced neural extensions such as Neural Hawkes Process [mei2017neural] [zuo2020transformer] and the Self-Attentive Hawkes Process [zhang2020self], which integrate RNNs or attention mechanisms into the intensity function. These models are particularly well suited for fine-grained timestamp prediction and have found applications in finance, healthcare, and user behavior modeling. However, TPPs typically assume a simple structure over events and focus solely on the temporal dimension, making them less suitable for capturing structured dependencies across users or networks.
Sequential recommendation.
A closely related application domain is sequential recommendation, where the goal is to predict the next item a user will interact with based on their historical behavior. Early methods employed Markov chains or matrix factorization over time-slided data [rendle2010factorizing], while more recent models use deep sequence encoders such as GRU4Rec [hidasi2015session], SASRec [kang2018self], and BERT4Rec [sun2019bert4rec], which apply Transformer-based architectures to item sequences. These models have shown strong performance by capturing user preferences over time. However, they typically model each user independently and do not account for interactions among users.
D.2 Graph models.
In parallel, there has been significant progress in graph-based modeling of user interaction, especially through Graph Neural Networks (GNN). While static graphs lack explicit timestamps and do not capture the order of interactions, they can still represent temporal data through careful graph construction, such as building graphs over specific time windows or pruning outdated edges [zhao2025gigl]. This setup allows for framing personal and relational event prediction tasks as link prediction (e.g. forecasting probability of user-item interactions) or node classification (e.g. fraud detection). Early GNNs like GCN [kipf2017semi] introduced neighborhood aggregation but were limited by their transductive nature and the requirement of full graph knowledge during training. GraphSAGE [hamilton2017inductive] addressed these issues by introducing inductive learning via stochastic sampling, while PinSAGE [ying2018graph] scaled GNNs to billions of nodes via relaxing memory constraints. GAT [velivckovic2018graph] incorporated attention mechanisms and later HGT [hu2020heterogeneous] extends GAT to heterogeneous graphs. These developments have established GNNs as a powerful tools for relational modeling [gao2023survey].
Temporal graph.
As described in gastinger2024tgb, temporal graph methods fall into two broad categories: discrete-time and continuous-time. Discrete-time methods exist for both homogeneous [han2020graph] and heterogeneous datasets [yin2019dhne, xue2021modeling, bian2019network]. Continuous-time methods arguably preserve more information and can be converted into discrete graphs, but the reverse is not possible [kazemi2020representation]. TGN [tgn_icml_grl2020] introduces a general framework for modeling continuous-time dynamic graphs, categorizing DyRep [trivedi2019dyrep] as a special case, followed by several other models such as DyGFormer [yu2023towards], NAT [luo2022neighborhood], TNCN [zhang2406efficient], and CTAN [gravina2024long]. Other temporal graph models such as HTGN-BTW [yue2022htgn] and STHN [li2023simplifying] propose different extensions of TGN to handle heterogeneous data. Beyond the standard temporal graphs, several methods have also been proposed for modeling temporal knowledge graphs [li2021temporal, jin2019recurrent, li2022complex].
D.3 Benchmark datasets.
Several benchmark efforts have been proposed across related areas. The temporal graph benchmarks include the TGB [huang2023temporal], its heterogeneous and knowledge graph extension (TGB 2.0) [gastinger2024tgb], and TGB-Seq [yi2025tgb], which include a more complex sequence of edge dynamics. For static graph learning, OGB [hu2020open] and its large-scale extension OGB-LSC [hu2021ogb] provide widely used benchmarks. In the recommendation domain, large-scale user-item interaction datasets have been released through benchmarks such as MIND [wu2020mind], TenRec [yuan2022tenrec], NineRec [zhang2024ninerec], and BARS [zhu2022bars]. For event sequence modeling and temporal point processes (TPP), recent benchmarks include EBES [osin2024ebes], EasyTPP [xue2023easytpp], and HOTPP [karpukhin2024hotpp].
The closest dataset or benchmark work from our paper is the temporal graph benchmark (TGB) by huang2023temporal. It contains several datasets that model user behaviors. These datasets can be roughly divided into two categories: (1) user-to-user interaction datasets, such as tgbl-coin, tgbl-comment, and tgbn-trade; (2) user-to-item bipartite interaction datasets, such as tgbl-wiki, tgbl-review, tgbn-genre, tgbn-reddit, and tgbn-token. The first category of user-to-user interaction datasets are similar to the relational event part of our datasets; however our datasets also contain personal event sequences, in addition to relational events, which are not present in the TGB dataset.
The second category of TGB datasets are bipartite temporal graphs. In our PRES formulation, the interaction of a user to an item can be encoded as a personal event, where an item is represented as an event or a token. However, the personal event abstraction in the PRES formulation can also encode other types of events. An example is illustrated in Figure 1, where User A has both user-item events (product views) and other types of events such as ‘Add to cart’ and ‘Purchase’. Our formulation also contains relational events that model use-to-user interactions.
In addition, formulating an item as a personal event instead of a node enables the flexibility of encoding items that have hierarchical information such as Geohash in our Brightkite dataset. Each character in the geohash encodes increasingly detailed location information. When we encode an 8-letter geohash as a node, we lose the hierarchical information encoded in the geohash. In contrast, if we are not forced to represent an event as a node, we have more flexibility to encode the hierarchical structure of the geohash. For example, in a sequence model, one could tokenize the event freely. A single event could be encoded into multiple tokens.
In terms of the size, the TGB dataset ranges from a small size of 255 node graph (tgbn-trade) to nearly a million node graph (tgbl-comment). Our PRES dataset also ranges from a small-to-medium size of 58 thousand users (pres-brightkite) to a relatively large dataset of pres-github with 3.6 million users. In terms of the number of events (or number of edges in TGB dataset) our PRES datasets are comparable with TGB datasets, and in some cases larger than TGB datasets. The number of edges in TGB datasets range from around 150 thousands edges (tgbl-wiki) to 44 million edges (tgbl-comment); whereas the number of events in our datasets range from 1.5 million (pres-amazon-clothing) to more than 100 million events (pres-github).
D.4 Other research on graph and sequence.
Several studies have been conducted on different settings on temporal and structural dynamics. Some models focus on modeling graph and time series data using spatio-temporal graph [yu2017spatio, yu2020spatio, bai2020adaptive, wu2020connecting]. Other model combine graph model output and sequence model output in various application areas [hai2024study, shi2019two, mao2021molecular]. A recent study focus on tokenizing graph and applying transformers or state space models (SSMs) for graph learning [zhang2001graph, wang2024graph, behrouz2024graph, huang2024can, ding2024dygmamba, li2024dyg, behrouz2024best]. Another studies works on incorporating knowledge graph into language model. [liu2020k, yasunaga2021qa, huang2020knowledge], as well as performing graph-augmented language generation [peng2024graph, zhu2025graph].
Appendix E Experiment Details
E.1 Hyperparameters
Personal event prediction task.
In Table 8, we present the hyperparameters used during the training of various models for personal event prediction tasks. We use the following notations: Emb Dim denotes the dimensionality of token embeddings; Heads is the number of attention heads; Layers refers to the number of hidden layers; Channels indicates the number of hidden channels per layer in GAT and GCN models; Max Examples is the maximum number of training samples generated per user; Num Neg Samples represents the number of negative samples for each (positive) sample; and Num Neighbors is the number of neighbors sampled per layer for GNN models. Additionally, due to GPU memory limitations, we reduce the embedding dimensions for the TGN and DyRep models to 16 and 32, respectively, for the pres-brightkite and pres-gowalla datasets, and to 32 and 64 for pres-amazon-clothing and pres-amazon-electronics.
Table 8: Hyperparameter configurations for personal event prediction tasks
Model Name Learning Rate Batch Size Epochs Emb Dim Heads Layers Channels Max Events Max Examples Num Neg Samples Num Neighbors
BERT 3e-4 1024 10 64 4 4 100 50 10
BERT-n2v-p 3e-4 1024 10 64 4 4 100 50 10
BERT-n2v-i 3e-4 1024 10 64 4 4 100 50 10
GCN 1e-3 1024 10 128 2 128 100 5 10
GCN_PR 1e-3 1024 10 128 2 128 100 5 10
GAT 1e-3 1024 10 128 2 2 64 100 5 10
GAT_PR 1e-3 1024 10 128 2 2 64 100 5 10
TGN 1e-3 4096 10 16/32 100 5 10
TGN_PR 1e-3 4096 10 16/32 100 5 10
DyRep 1e-3 4096 10 32/64 100 5 10
DyRep_PR 1e-3 4096 10 32/64 100 5 10
Relational event prediction task.
In Table 9, we present the hyperparameters used across all models for relational event prediction tasks. Due to memory and time constraints, batch size, number of epochs, and embedding dimensions were adjusted per dataset. All datasets used a batch size of 4096, except for pres-github, which used 512. The number of training epochs was set to 5 for pres-github, 20 for pres-gowalla and pres-amazon-electronics, 100 for pres-amazon-clothing, and 1000 for pres-brightkite. The model checkpoint with the best validation MRR was saved and used for testing. As shown in our results, TGN, DyRep and TNCN could not be run on pres-github. For the remaining datasets, the embedding dimension for TGN and DyRep was 128, except for pres-gowalla, which used 64 to avoid GPU out-of-memory errors. TNCN used ‘NCN_mode‘ of 1, an embedding dimension of 64 and a smaller batch size (1024) for all datasets.
Table 9: Hyperparameter configurations for relational event prediction tasks
Model Name Learning Rate Batch Size Epochs Emb Dim Heads Layers Channels Num Neg Samples Num Neighbors
GCN 1e-3 512/4096 5-1000 128 2 128 5 10
GCN_PR 1e-3 512/4096 5-1000 128 2 128 5 10
GCN+S 1e-3 512/4096 5-1000 128 2 128 5 10
GAT 1e-3 512/4096 5-1000 128 2 2 128 5 10
GAT_PR 1e-3 512/4096 5-1000 128 2 2 128 5 10
GAT+S 1e-3 512/4096 5-1000 128 2 2 128 5 10
TConv 1e-3 512/4096 5-1000 128 2 2 128 5 10
TConv_PR 1e-3 512/4096 5-1000 128 2 2 128 5 10
TConv+S 1e-3 512/4096 5-1000 128 2 2 128 5 10
TGN 1e-3 4096 20-1000 64/128 128 5 10
DyRep 1e-3 4096 20-1000 64/128 128 5 10
TNCN 1e-3 1024 20-1000 64 128 5 10
Table 10: Computational Time (in hours) for Different Models and Datasets
Method Time (h)
amazon-clothing amazon-electronics brightkite gowalla github
Relational event prediction tasks
[…]
Personal event prediction tasks
[…]
E.2 Computing Resources
We conducted all experiments on a server equipped with 8 NVIDIA Ampere A10G GPUs (24 GB each), 16 CPU cores, and a RAM upper limit of 512 GB. To fully leverage all resources, we parallelized the training runs so that each experiment used a single GPU. Each experiment is designed to be run on a single-GPU machine. Table 10 summarizes the average training time (in hours) and standard deviation for each model across five datasets, categorized by task type. For relational event prediction tasks, lightweight GCN and GAT variants exhibit minimal computational overhead, with training times generally under one hour except on the GitHub dataset. In contrast, temporally expressive models such as TGN and DyRep incur significantly higher costs, especially on large-scale datasets like Gowalla. In personal event prediction tasks, training times increase across the board, with most models exceeding 7 hours on larger datasets, again highlighting the computational demands of modeling fine-grained temporal dynamics.
Figure 4: Comparison of relational event predictions across different datasets.
Appendix F Additional Experimental Results
Experiment Figures
In Figures 4 and 5, we present the results from the main paper in a more visual format to facilitate comparison across methods. In the relational event prediction tasks, across all datasets and metrics, static GNNs augmented with personal event sequence embeddings (GCN+S, GAT+S, and especially TConv+S) consistently perform well, achieving the best or second-best results. This highlights the benefit of integrating both personal and relational signals. For temporal graph methods, the TGN and DyRep under-perform in most of datasets and most metrics. TNCN perform well on amazon datasets on MRR and Hits@k with lower $k$, but under-perform on other metrics or other datasets. For personal event prediction tasks, BERT+n2v-i offers slight improvements over regular BERT. In particular, BERT-based models exhibit competitive performance in some cases, most prominently on the Brightkite dataset, where they outperform GNN-based counterparts at MMR and lower hit rate thresholds such as Hits@3, Hits@5, and Hits@10.
Figure 5: Comparison of personal event predictions across different datasets.
Additional analysis
One of the main takeaways of the paper is that models leveraging both personal and relational events outperform those using only one type in either relational or personal event prediction tasks. For example, in relational event prediction, the "+S" models (GCN+S, GAT+S, and TConv+S) incorporate sequence embeddings of personal event data into the relational event graph, boosting performance over models that rely solely on relational events (GCN, GAT, or TConv). This highlights the need for models that jointly account for both signals.
We then compare the "+S" strategy to the "_RP" models, which convert each unique personal event into a node and add it to the user-to-user graph used by the GCN model, creating edges between users and their personal event nodes. In most cases, "+S" models outperform "_RP" models, with very few exceptions. In some datasets, such as pres-brightkite, pres-gowalla, and pres-github, adding personal event nodes to the graph even reduces performance. These results may be explained as follows:
  • When we exclude personal events (standard GCN, GAT, and TConv), the model is still able to extract and learn some predictive information from the user-to-user relational events alone to some degree.
  • When we include personal events as nodes (GCN_RP, GAT_RP, TConv_RP), this adds more noise than benefit to the system, as the number of personal event nodes is much larger than the number of user nodes. As a result, model performance decreases.
  • When we encode personal events as sequence embeddings (GCN+S, GAT+S, and TConv+S), this produces meaningful features without introducing excessive noise. The models are able to capture additional signals from these personal event embeddings, leading to performance improvements.
  • In addition, when modeling personal events using a sequence model (BERT) in the "+S" strategy, we retain the hierarchical information of the personal events (such as geohash check-in events in pres-brightkite and pres-gowalla). In contrast, when we convert personal events into nodes as in the "_RP" strategy, we lose the hierarchical information present in the events.
However, this pattern does not always generalize to every dataset, as we see in pres-amazon-clothing, where "_RP" models perform relatively well on MRR and Hits@5, but not on Hits@k metrics with larger $k$. This suggests that in this dataset, personal event nodes may not merely act as noise in the graphs. Instead, they help the model improve precision on the top candidates (i.e., fewer but more accurate suggestions), at the expense of lower recall coverage for larger $k$. In addition, the encoded hierarchical information in the product-rating nodes may be less important in this dataset. This observation may influence the architecture design of future models aiming to leverage both personal and relational event signals.
Appendix G Broader Impacts
Broader impact
The datasets and prediction tasks we release may support future research on user event modeling, particularly in settings that involve both personal and relational events. Researchers can build models on top of these resources and evaluate them in a consistent way. This can help accelerate empirical progress and facilitate more comparable results. This has potential impact in a range of industry applications where modeling user behavior is critical, such as recommendation, fraud detection, and user interaction analysis.
Potential negative impact
The datasets we release may not cover all use cases of user event modeling, and may reflect only a subset of real-world scenarios. This could introduce bias in model development or evaluation, especially if models are tuned specifically for the structure or properties of our datasets. As a result, there is a risk that future methods may overfit to our datasets and generalize less effectively to other domains or applications.